by seeing a child on a railway track the driver applies brakes the speed of train is reduced from 90 km/h to 36km/h in 5s then train retards by


1
Expert's answer
2021-10-19T15:07:28-0400

The retardation is


"a=\frac{v_f-v_i}{t}=\frac{36-90}{5}=-10.8\text{ km/h/s}."

If we convers the speed to m/s, we have


"a=-10.8·1000/3600=-3\text{ m/s}^2."